OpenStudy (anonymous):

find the point on the graph of f(x)=x^2 that is closest to (4,6)

OpenStudy (jmark):

Let we are looking for the point (x,y) on f(x)=x^2 that is closest to (4,6) =>y = x^2 The slope of the tangent at that point is =>y' = 2x The slope of the line between (x,y) and (4,6) = -1/y' =>-1/2x = (y – 6)/(x-4) =>-(x-4)=2xy-12x =>2x^3-11x-4=0 =>x =2.5094 =>x=-2.1363 and x=-0.37308 So, (2.5094,6.29) is the point on f(x) closest to (4,6). To double check it graphically, it appears correct, as the smallest circle centered at (4,6) that touches f(x) seems to be at x=2.5094 Know more Sinusoidal Graph http://goo.gl/MOJm9U